OpenStudy (anonymous):

 the equation 6x − 1 = 11, which order of operations completely solves for x

OpenStudy (anonymous):

6x - 1 = 11 6x = 11 + 1 \[x = \frac{ 11 + 1 }{ 6 }\]

OpenStudy (anonymous):

x=2

OpenStudy (texaschic101):

order of operations ?.....first you add 1 to both sides, then you divide both sides by 6. Just like Hamoody did